# TinkerXR: In-Situ, Reality-Aware CAD and 3D Printing Interface for Novices

### v3

@@ 6.3. Qualitative Feedback
Participants consistently praised the ability to design within real environments, emphasizing the value of TinkerXR’s AR capabilities in visualizing objects in situ. For instance, P1 shared, “With TinkerXR, I could see how my design would actually look in the space it was meant for.” P4 highlighted how designing anywhere — including outdoors — opened new possibilities, such as creating projects in natural settings. The immersive experience was widely appreciated, with features like depth occlusion enhancing realism. As P10 noted, “The occlusion made everything feel so much more immersive, like the objects were truly part of the room.” The engaging and enjoyable nature of the system was also mentioned, with P3 stating, “Using the system was genuinely fun and engaging.”
Intuitive and Natural Controls
The hand-based interactions and intuitive controls stood out as significant strengths. P6 described them as “human-nature,” explaining, “Locking axes was so simple, just a left-hand finger pinch and moving along the axis I wanted.” Similarly, the snap-to-grid feature was highly valued for its precision, with P2 commenting, “The snap-to-grid feature was a game-changer for precise alignments.” These intuitive mechanisms allowed participants to focus on creativity without being hindered by complex operations.
@@
While the controls were intuitive, several participants expressed concerns about the physical demands of prolonged use. P1 mentioned, “The physical demand was rather tiring,” and P2 noted, “It was a little tiring to have my arm raised for every action.” These observations suggest that extended use could lead to fatigue, especially when designing complex models.
@@
TinkerXR was widely commended for its accessibility, even among novice users. P6 shared, “As someone completely new to AR, 3D modeling, and 3D printing, I was surprised by how easy it was to learn.” This ease of use motivated participants to envision future applications, such as designing custom objects for their homes or gardens. P4 expressed excitement about using the system for personal projects, stating, “Even in my garden, I am already thinking about designing a birdhouse and some tree decorations.” Most participants indicated they would continue using the tool and recommend it to others, with P8 emphasizing its convenience: “Starting the print directly from the headset saves materials because I don’t have to redo prints.”
@@
Participants identified areas where additional features could enhance the workflow. Several requested better tools for manipulating small objects, with P7 suggesting a “magnifying glass tool for designing small objects.” Building on the AR ruler feature, on-screen measurements were another commonly desired functionality, as P1 noted, “Length values of the edges appearing next to the objects, just like in Tinkercad, would help.” Other recommendations included importing external 3D models and freely adjusting grid sizes. P4 also proposed more direct interaction methods, such as “grabbing objects directly with our hands.”
Overall, the feedback highlights TinkerXR’s strengths in providing an immersive, intuitive, and accessible design experience, especially for novices. However, challenges like solving the problem of physical fatigue on extended usage should be solved in the future. Addressing functionalities like small object manipulation and hand-grabbing the objects would also further enhance its usability.
@@ 7. Discussion
Our tool was designed with novices in mind, providing an intuitive interface that simplifies both the design and fabrication of models. However, we envision that TinkerXR can also be beneficial to experts. Practitioners and professional designers can use it for rapid prototyping, concept development, and interactive presentations. The ease of use and flexibility of our system can streamline workflows, allowing experts to focus on more complex design tasks without being bogged down by technical logistics. In the future, we believe that our system can be adopted by educational institutions for teaching design principles, by industries for quick mock-ups or collaborative sessions, and by hobbyists seeking to explore 3D modeling and fabrication.
@@
In Section 5, we demonstrated our tool’s usefulness through various design examples, each showing a different use case. These objects featured distinct geometries, ranging from small details to larger-scale structures. For instance, the candy bowl model in Figure 14d is made of 32 curved lines, each made of 8 cuboids, adding up to a total of 256 basic building blocks (if none of the blocks are combined) and 2048 vertices. Our system supports up to 300 basic building blocks and a total number of vertices of 20,000 in real time. However, we observed a noticeable lag when exceeding these numbers, especially in combination operations. We anticipate this limitation can be addressed with future headsets offering higher processing power. Alternatively, more powerful headsets, e.g., Varjo XR-4141414https://varjo.com/products/xr-4/, Accessed: 4.08.2024, which is connected to a desktop computer, can support compute-intensive tasks.
While the system can create a variety of complex geometries, certain tasks remain challenging due to the lack of more advanced tools such as extrusion of 3D geometries from 2D blueprints (e.g., sketched shapes, text), parametric curves, lofting, filleting, freeform design tools similar to those in Autodesk 3ds Max [^15] https://www.autodesk.com/products/3ds-max, Accessed: 4.08.2024, or a collaborative structure allowing multiple users to work on the same project similar to Gravity Sketch or adding physical engagement with local users and their surroundings similar to HoloBots (Ihara et al., 2023). These features were not implemented, as our system targets novices and prioritizes simplicity. While it is arguable that adding such features inherently increases complexity, our design choices focus on maintaining an intuitive user experience.
We only implemented the features that were required for realizing the artifacts illustrated in Section 5 and objects with similar structural complexity, which were inspired by Tinkercad repositories and related work. In the future, because we open-source our work, we believe that the community can add more features to enhance its capabilities. By fostering a collaborative development environment, we hope to see our system evolve to meet the needs of users of every proficiency, ultimately broadening its applicability and usability.
@@
In this context, Roumen et al. (Baudisch et al., 2019) provide inspiration for adapting 3D design for laser cutting, although it is not a general-purpose solution. Exploring ways to integrate such techniques into TinkerXR could enable users to understand fabrication-specific constraints in real time, which we consider an exciting avenue for future work.
@@
MR headsets can be bulky; however, in our user study, participants did not report major problems due to the form factor. We used the Quest 3 because it is lightweight, making it easy to use. In the future, we imagine AR technology becoming more ubiquitous while offering a more streamlined experience with the advent of AR glasses, such as Meta Orion [^16] https://about.meta.com/realitylabs/orion/, Accessed: 20.01.2025. Future devices are likely to be more compact, resembling ordinary eyewear, thus enhancing user comfort and accessibility (Dogan, 2024). These advancements could lead to broader adoption and more immersive systems.
@@
While we are constrained by Meta’s hand gesture APIs, which are known to improve over time [^17] https://www.techradar.com/computing/virtual-reality-augmented-reality/your-meta-quest-3-is-getting-a-hand-tracking-upgrade-that-could-unlock-foot-tracking, Accessed: 5.08.2024, there are still issues such as improper functionality in dimly lit areas or limited precision of tracking compared to controllers or the system being exclusive to Meta headsets.
@@ 8. Conclusion
In this paper, we presented TinkerXR, a new 3D modeling and fabrication tool through AR, designed for novices. Our approach lets users design artifacts with a simple primitive geometry manipulation CSG design logic, with intuitive hand gesture control schemes, on the intended space of the artifact, and with seamless integration with the 3D printing process. We showed the various features present in TinkerXR from selecting the workspace to design and fabrication. Using these features, we presented a wide variety of creations for different scenarios, each varying in complexity and practical applications. A user study revealed that the system was effective for novices, and more immersive than Tinkercad, while offering a comparable level of functionality and simplicity. We hope that our project and its open-source codebase will stimulate further exploration within communities engaged in HCI, design, and fabrication.

### v4

@@ 6.3. Qualitative Feedback
Participants consistently praised the ability to design within real environments, emphasizing the value of TinkerXR’s AR capabilities in visualizing objects in situ. For instance, P1 shared, “With TinkerXR, I could see how my design would actually look in the space it was meant for.” P4 highlighted how designing anywhere — including outdoors — opened new possibilities, such as creating projects in natural settings. The immersive experience was widely appreciated, with features like depth occlusion enhancing realism. As P10 noted, “The occlusion made everything feel so much more immersive, like the objects were truly part of the room.” The engaging and enjoyable nature of the system was also mentioned, with P3 stating, “Using the system was genuinely fun and engaging. I felt immersed.”
Natural Controls
The hand-based interactions stood out as significant strengths. P6 described them as “human-nature,” explaining, “Locking axes was so simple, just a (weak-hand) finger pinch and moving your hand on a direction.” Similarly, the snap-to-grid feature was highly valued for its precision, with P2 commenting, “snap-to-grid was a game changer for precise alignments.” and P3 commenting, ”Snapping was very nice since it helped against my hand shaking on air.” These mechanisms allowed participants to focus on creativity without being hindered by complex control schemes.
@@
While the controls felt natural, several participants expressed concerns about the physical demands of prolonged use. P1 mentioned, “The physical demand was rather tiring,” and P2 noted, “It was a little tiring to have my arm raised for every action.” These observations suggest need of improvements on reducing the physical effort required to perform operations in TinkerXR. It can be argued that with the current system, extended use could lead to fatigue, especially when designing complex models.
@@
TinkerXR was widely commended for its ease of use and intuitiveness, even among novice users. P6 shared, “As someone completely new to AR, 3D modeling, and 3D printing, I was surprised by how easy it was to learn.” This ease of use motivated participants to envision future applications, such as designing custom objects for their homes or gardens. P4 expressed excitement about using the system for personal projects, stating, “Even in my garden, I am already thinking about designing a birdhouse and some tree decorations.” Most participants indicated they would continue using the tool and recommend it to others, with P8 emphasizing a room for a collaborative design directon: “This was a fun experience overall. I would recommend people to try it out. I wish I could design simultaneously with others in the same room too – that would have been awesome.”
@@
On top of collaboration features, participants identified areas where additional design features could enhance the workflow. Several requested better tools for manipulating small objects, with P7 suggesting a “magnifying glass tool for designing small objects and P1 mentioning ”Small objects was hard to select and manipulate sometimes because I needed to be very precise with my hand.” Building on the AR ruler feature, on-screen measurements were another commonly desired functionality, as P1 noted, “Length values of the edges appearing next to the objects, just like in Tinkercad, would help.” Other recommendations included utilizing generative design and speech recognition to further simplify the design process. P4 also proposed more direct interaction methods, such as “grabbing objects directly with our hands (palm grab).”
Overall, the feedback highlights TinkerXR’s strengths in providing an immersive and accessible design experience, especially for novices. However, challenges like solving the problem of high physical demand should be solved in the future. Addressing the lack of precision while manipulating small objects would further enhance the system’s usability.
@@ 7. Discussion
Our tool was designed with novices in mind, with the aim of providing an intuitive interface that simplifies both the design and fabrication of models. However, we envision that TinkerXR can also be beneficial to experts. Practitioners and professional designers can use it for rapid prototyping, concept development, and interactive presentations. The ease of use and flexibility of our system can streamline workflows, allowing experts to focus on more complex design tasks without being bogged down by technical logistics. In the future, we believe that our system can be adopted by educational institutions for teaching design principles, by industries for quick mock-ups or collaborative sessions, and by hobbyists seeking to explore 3D modeling and fabrication.
@@
In Section 5, we demonstrated our tool’s usefulness through various design examples, each showing a different use case. These objects featured distinct geometries, ranging from small details to larger-scale structures. For instance, the candy bowl model in Figure 11d is made of 32 curved lines, each made of 8 cuboids, adding up to a total of 256 basic building blocks (if none of the blocks are combined) and 2048 vertices. Our system supports up to 300 basic building blocks and a total number of vertices of 20,000 in real time. However, we observed a noticeable lag when exceeding these numbers, especially in combination operations. We anticipate this limitation can be addressed with future headsets offering higher processing power. Alternatively, more powerful headsets, e.g., Varjo XR-4151515https://varjo.com/products/xr-4/, Accessed: 4.08.2024, which is connected to a desktop computer, can support compute-intensive tasks.
While the system can create a variety of complex geometries, certain tasks remain challenging due to the lack of more advanced tools such as extrusion of 3D geometries from 2D blueprints (e.g., sketched shapes, text), parametric curves, lofting, filleting, freeform design tools similar to those in Autodesk 3ds Max [^16] https://www.autodesk.com/products/3ds-max, Accessed: 4.08.2024, or a collaborative structure allowing multiple users to work on the same project similar to Gravity Sketch or adding physical engagement with local users and their surroundings similar to HoloBots (Ihara et al., 2023). While it is arguable that adding such features may increase complexity, their addition would elaborate the design capabilities, and should be considered as a future improvement.
We only implemented the features that were required for realizing the artifacts illustrated in Section 5 and objects with similar structural complexity, which were inspired by Tinkercad repositories and related work. In the future, because we open-source our work, we believe that the community can add more features to enhance its capabilities. We anticipate future research to also support more advanced workflows, such as multi-material or -texture prints (Ozdemir et al., 2024). By collaborating with the community, we hope to see our system evolve to meet the needs of users of all proficiency levels.
@@
In this context, kyub (Baudisch et al., 2019) provides inspiration for adapting 3D design for laser cutting. Exploring ways to integrate such techniques into TinkerXR could enable users to understand fabrication-specific constraints in real time, which we consider an exciting avenue for future work.
@@
MR headsets can be bulky; however, in our user study, participants did not report major problems due to the form factor. We used the Quest 3 because it is lightweight, making it easy to use. In the future, we imagine AR technology becoming more ubiquitous while offering a more streamlined experience with the advent of AR glasses, such as Meta Orion [^17] https://about.meta.com/realitylabs/orion/, Accessed: 20.01.2025. Future devices are likely to be more compact, resembling ordinary eyewear, thus enhancing user comfort and accessibility (Dogan, 2024). These advancements could lead to broader adoption and more immersive systems (Feick et al., 2025). Currently, TinkerXR is implemented exclusively for the Meta Quest ecosystem. However, its core architecture, built in Unity, is fundamentally portable. Expanding the system to other platforms with robust hand tracking and spatial mapping capabilities, such as the Apple Vision Pro or Varjo headsets, represents a clear path for future development and would broaden the system’s accessibility.
@@
While we are constrained by Meta’s hand gesture APIs, which are known to improve over time [^18] https://www.techradar.com/computing/virtual-reality-augmented-reality/your-meta-quest-3-is-getting-a-hand-tracking-upgrade-that-could-unlock-foot-tracking, Accessed: 5.08.2024, there are still issues such as improper functionality in dimly lit areas or minor precision errors, especially during fine-scale manipulation. Our study showed that participants effectively used the snap-to-grid feature to mitigate this. To build on these strategies, future work could explore several paradigms to enhance precision further. For example, implementing a virtual magnifying glass tool, as suggested by user feedback, would offer a stable, magnified view for detailed edits. Similarly, enabling users to fluidly work at a larger, more manageable scale before uniformly shrinking their design could simplify the manipulation of intricate components. Finally, extending the system to support dynamic object-to-object snapping (e.g., aligning vertices, edges, and faces) would bring its precision capabilities closer to traditional desktop controls.
Utilization of Generative AI
Our findings underscore the effectiveness of CSG for novices, yet the 3D Geometry Complexity discussion highlights its inherent limitations in creating freeform or highly intricate models. A compelling future research direction is the integration of generative AI to address this expressivity gap. In such a hybrid workflow, a user could leverage TinkerXR’s direct manipulation capabilities to sculpt a low-fidelity ”scaffold” of their design. This user-authored geometry, serving as a strong structural prior, could then guide state-of-the-art generative models to refine and detail the final object. Recent developer access to Quest 3’s cameras [^19] https://github.com/oculus-samples/Unity-PassthroughCameraApiSamples, Accessed: 8.09.2025 also opens the opportunity for these models to use environmental imagery to ensure the final design is aesthetically consistent with the user’s room. This approach would preserve user agency where the foundational design intent remains user-driven and AI acts as a tool for complex refinement rather than the sole creator (Drago et al., 2025). Investigating this interplay between direct manipulation and AI-driven detailing, and its effect on novice users’ sense of authorship presents an interesting research question for the next generation of XR design tools.
@@ 8. Conclusion
In this paper, we presented TinkerXR, a new 3D modeling and fabrication tool through AR, designed for novices. Our approach lets users design artifacts with a simple primitive geometry manipulation CSG design logic, with intuitive hand gesture control schemes, on the intended space of the artifact, and with seamless integration with the 3D printing process. We showed the various features present in TinkerXR from selecting the workspace to design and fabrication. Using these features, we presented a wide variety of creations for different scenarios, each varying in complexity and practical applications. A user study revealed that the system was effective for novices, and more immersive than Tinkercad, while offering a comparable level of functionality and simplicity. We hope that our open-source project will stimulate further exploration within communities engaged in HCI, design, and fabrication.